# Neural Concept Verifier: Scaling Prover-Verifier Games via Concept Encodings

### v2

@@ Abstract
While Prover-Verifier Games (PVGs) offer a promising path toward verifiability in nonlinear classification models, they have not yet been applied to complex inputs such as high-dimensional images. Conversely, Concept Bottleneck Models (CBMs) effectively translate such data into interpretable concepts but are limited by their reliance on low-capacity linear predictors. In this work, we introduce the Neural Concept Verifier (NCV), a unified framework combining PVGs with concept encodings for interpretable, nonlinear classification in high-dimensional settings. NCV achieves this by utilizing recent minimally supervised concept discovery models to extract structured concept encodings from raw inputs. A prover then selects a subset of these encodings, which a verifier—implemented as a nonlinear predictor—uses exclusively for decision-making. Our evaluations show that NCV outperforms CBM and pixel-based PVG classifier baselines on high-dimensional, logically complex datasets and also helps mitigate shortcut behavior. Overall, we demonstrate NCV as a promising step toward performative, verifiable AI.
@@ 1 Introduction
Deep learning achieves remarkable predictive performance at the expense of interpretability and trustworthiness (Rudin,, 2019). However, particularly in high-stakes applications, it is critical that models provide verifiable justifications for their decisions (Irving et al.,, 2018; Fok and Weld,, 2023). Prover-Verifier Games (PVGs), introduced by Anil et al., (2021), formalize such justifications via a game-theoretic approach: a prover provides evidence to convince a verifier, who accepts only verifiable proofs. A prominent instantiation of PVGs is the Merlin-Arthur Classifier (Wäldchen et al.,, 2024), i.e., a classifier guided by cooperative and adversarial provers, offering formal interpretability guarantees through information-theoretic bounds. However, the Merlin-Arthur framework faces significant scalability challenges when applied to high-dimensional real-world data, as explanations based on raw pixels are both computationally difficult to optimize and offer limited human understandability (cf. Figure  1; Wäldchen et al., (2024)).
@@
Figure 2: Overview of the Neural Concept Verifier (NCV). The input image is first processed by a concept extractor to produce symbolic concept encodings. A Prover–Verifier Game is then played over these encodings: a cooperative prover selects a sparse concept subset supporting the true class, while an adversarial prover selects misleading concepts. Finally, the verifier makes a prediction based only on these selected concepts, ensuring faithful and robust classification.
@@
Concurrently, Concept Bottleneck Models (CBMs) have emerged as a powerful framework for interpretable machine learning, structuring predictions through intermediate interpretable concept encodings (Koh et al.,, 2020; Stammer et al.,, 2021). Despite their advantages, CBMs typically employ linear classifiers on top of concept encoding layers, thereby potentially restricting expressivity and failing on tasks requiring nonlinear interactions among concepts (e.g., XOR problems, counting or permutation invariance) (Mahinpei et al.,, 2021; Kimura et al.,, 2024; Lee et al.,, 2019).
@@
In this work, we bridge these two complementary directions by introducing the Neural Concept Verifier (NCV), a novel framework integrating concept-based representations into PVGs in the form of Merlin-Arthur Classifiers. NCV shifts the prover–verifier interaction from the image level to a structured, symbolic concept level, overcoming both the scalability limitations encountered by Merlin-Arthur Classifiers in high-dimensional settings and the expressivity constraints inherent to linear Concept Bottleneck Models. Through extensive evaluations on controlled synthetic benchmarks such as CLEVR-Hans, as well as large-scale real-world datasets including CIFAR-100 and ImageNet-1k, we demonstrate that NCV successfully scales PVGs to complex, high-dimensional classification tasks. Additionally, NCV enables verifiable, performant nonlinear classifiers on top of concept extractors, effectively narrowing the interpretability–accuracy gap present in standard, linear CBMs. Lastly, our framework improves robustness to shortcut learning, thereby enhancing the generalizability and trustworthiness of predictions, particularly crucial in high-stakes applications.
@@
Our contributions can be summarised as: (i) We propose Neural Concept Verifier (NCV), a framework combining concept-based models with Prover–Verifier Games (PVGs). (ii) NCV scales PVGs to high-dimensional image data by operating on compact concept encodings. (iii) It enables expressive yet interpretable classification via sparse, nonlinear reasoning over concepts. (iv) We validate NCV on synthetic and real-world benchmarks, demonstrating strong accuracy and verifiability. (v) We highlight that NCV improves generalization under spurious correlations, indicating increased robustness.
The remainder of the paper is structured as follows. We begin with a review of related work, highlighting recent developments in the field. We then introduce our proposed Neural Concept Verifier framework, and present its formal description. This is followed by a comprehensive experimental evaluation that investigates key aspects of the framework. Finally, we discuss our findings and conclude the paper.
@@ 2 Background
Prover-Verifier Games. PVGs were introduced by Anil et al., (2021) as a game-theoretic framework to encourage learning agents to produce testable justifications through interactions between an untrusted prover and a trusted verifier. Their work showed that under suitable game settings, the verifier can learn robust decision rules even when the prover actively attempts to persuade it of arbitrary outputs. Wäldchen et al., (2024) extended this idea to the Merlin-Arthur Classifier (MAC), which provides formal interpretability guarantees by bounding the mutual information between the selected features and the ground-truth label. Recently, Kirchner et al., (2024) applied a PVG-inspired approach to improve legibility of Large Language Model (LLM) outputs and Amit et al., (2024) introduced self-proving models that leverage interactive proofs to formally verify the correctness of model outputs. PVG-style setups have also been explored in safety-focused learning protocols (Irving et al.,, 2018; Brown-Cohen et al.,, 2024; Głuch et al.,, 2024). These developments reflect a broader trend of utilising multi-agent learning (Pruthi et al.,, 2022; Schneider and Vlachos,, 2024; Du et al.,, 2024; Nair et al.,, 2023; Stammer et al., 2024a, ). Our work builds on this line of research by embedding PVGs into a concept-based classification framework, addressing the dimensionality bottleneck that limits PVGs in high-dimensional settings.
@@
Concept Representations for Interpretability. The introduction of Concept Bottleneck Models (CBMs) (Koh et al.,, 2020; Delfosse et al.,, 2024), but also concept-based explanations (Kim et al.,, 2018; Crabbé and van der Schaar,, 2022; Poeta et al.,, 2023; Lee et al.,, 2025) was an important moment in the growing interest in AI interpretability research. The particular appeal of CBMs lies in the promise of interpretable predictions and a controllable, structured interface for human interactions (Stammer et al.,, 2021). While initial CBMs relied on fully supervised concept annotations, subsequent research has relaxed this requirement by leveraging pretrained vision-language models like CLIP for concept extraction (Bhalla et al.,, 2024; Yang et al.,, 2023; Oikarinen et al.,, 2023; Panousis et al.,, 2024), or employing fully unsupervised concept discovery methods (Ghorbani et al.,, 2019; Stammer et al., 2024b, ; Schut et al.,, 2025; Sawada and Nakamura,, 2022). While previous work has focused on enhancing the concept bottleneck itself, by reducing supervision requirements, mitigating concept leakage (Havasi et al.,, 2022), or dynamically expanding the concept space (Shang et al.,, 2024), the classifier component of CBMs, has received little attention. We argue that while a linear layer offers interpretability, its limited expressiveness makes it unsuitable for handling complex, compositional, and object-centric tasks. Our proposed NCV tackles this limitation.
@@
Shortcut Learning. Independent of their interpretability, training deep models can often lead to unwanted artifacts and side effects. Shortcut learning describes the problem of models learning to rely on unwanted and unintended features to resolve a task (Geirhos et al.,, 2020). This is a common problem when training purely deep learning models (Lapuschkin et al.,, 2019; Schramowski et al.,, 2020) or neuro-symbolic models (Marconato et al.,, 2023; Bortolotti et al.,, 2025), and, if not taken care of, can lead to predictions being right for the wrong reasons (Ross et al.,, 2017). There have been various methods developed to tackle this problem, from careful dataset curation (Ahmed et al.,, 2021) to modified model training (Friedrich et al.,, 2023), cf. (Steinmann et al.,, 2024) for a comprehensive overview. However, these mitigation methods rely on several important assumptions about the training data and potential shortcuts to mitigate, and can affect model performance (Sagawa et al.,, 2019). While NCV is not specifically designed to mitigate shortcuts, we show that our setup can intrinsically mitigate the impact of shortcuts in the data.
@@ 3 Neural Concept Verifier (NCV)
In this section, we introduce the Neural Concept Verifier (NCV), a novel framework for training interpretable and verifiable nonlinear classifiers capable of handling high-dimensional inputs. NCV combines concept-based representations with the Prover-Verifier Game setup of MAC (Wäldchen et al.,, 2024), enabling models whose predictions are certified to rely on a sparse set of high-level concepts. This approach supports both model interpretability and predictive performance, making concept-based models more competitive with opaque architectures and at the same time extending the benefits of PVGs to high-dimensional data. NCV consists of two main components: (i) a minimally or weakly supervised concept extractor that transforms input data into interpretable concept encodings, and (ii) MAC that selects and verifies sparse concept-based inputs to make final predictions. An overview of the framework is shown in Figure 2. After providing background notations and a high-level overview of the NCV framework below, we provide detailed descriptions of each main component as well as training and inference details in the following.
@@ 3.1 Problem Setup and Notation
Let $\mathcal{X}\in\mathbb{R}^{N\times D}$ denote a dataset of $N$ inputs (e.g., images), each of dimension $D$, and $\mathcal{Y}\in\mathbb{R}^{N\times(K+1)}$ the corresponding labels for $K$ classes plus one additional rejection class. The rejection class allows the model to abstain from a decision when uncertain, which is important for enforcing interpretability guarantees in adversarial setups (Wäldchen et al.,, 2024). Our overall goal is to learn a model $f:\mathcal{X}\rightarrow\mathcal{Y}$ that maps an input $x\in\mathcal{X}$ to a prediction $\hat{y}\in\mathcal{Y}$, where the prediction is based on a small, interpretable subset of high-level concepts.
@@ 3.2 The NCV Framework
NCV next decomposes $f$ into the following components (cf. Figure  2 for an illustrative overview):
@@
    A pair of prover agents $M,\widehat{M}:\mathcal{C}\rightarrow\{0,1\}^{C}$ that produce sparse binary masks selecting $m$ concepts each. $M$ (Merlin, cooperative Prover) aims to help classification; $\widehat{M}$ (Morgana, adverserial Prover) aims to mislead.
@@
    A verifier (Arthur) $A:\mathbb{R}^{C}\rightarrow\mathcal{Y}$, which predicts a label based only on the masked concepts.
@@ 3.3 Concept Extraction
The concept extractor $g$ transforms input images into interpretable, high-level feature encodings. While NCV is compatible with any concept bottleneck or discovery model, we experiment with two instantiations in our evaluations:
CLIP-based Concept Extraction.
In one instantiation we leverage pretrained vision-language models such as CLIP (Radford et al.,, 2021). Given an image $x$ and a vocabulary of textual concept descriptions $T=\{t_{1},...,t_{C}\}$, we compute: $\mathbf{c}_{i}=\langle e_{\text{img}}(x),e_{\text{text}}(t_{i})\rangle\quad\forall i\in\{1,...,C\}$, where $e_{\text{img}}$ and $e_{\text{text}}$ are CLIP’s image and text encoders, respectively. The resulting concept encoding $\mathbf{c}\in\mathbb{R}^{C}$ measures the similarity of the image to each textual concept. Unlike sparse activation methods such as SpLiCE (Bhalla et al.,, 2024), which enforce sparsity directly in the concept activation space $\mathbf{c}$ to promote human interpretability, we shift this constraint to the level of concept usage. Thus, in NCV, we maintain a dense and expressive concept space but enforce sparsity at the downstream classifier level, where only a small subset of concepts is selected for each decision. This results in task-driven sparsity that emerges naturally through the prover–verifier interaction.
Unsupervised Concept Extraction via NCB.
With our second instantiation we demonstrate that NCV is also compatible with unsupervised concept discovery modules. For this, we instantiate $g$ using the Neural Concept Binder (NCB) (Stammer et al., 2024b, ) that infers discrete, symbolic concepts from raw images $\mathcal{X}$, without any supervision. It first produces soft, continuous object-centric encodings $z=g_{\text{soft}}(x)$ using slot-based attention and factor binding. These are then discretized via retrieval-based clustering, resulting in hard, symbolic concept encodings $\mathbf{c}=g_{\text{hard}}(z)$, leading to $\mathbf{c}=(g_{\text{hard}}(g_{\text{soft}}(x))\in\mathbb{R}^{C}$, where each entry corresponds to a learned visual concept. Importantly, NCB’s concept encodings (in comparison to those extracted via CLIP) are object-centric, i.e., represent concepts per object in an image. Since NCB does not rely on textual prompts or class labels, its concept space may include noisy or redundant concepts. However, the sparse concept selection in NCV mitigates this by identifying only the most relevant concepts for each prediction.
The goal of investigating these two concept extractor instantiations is to demonstrate NCV’s generality and provide a good testbed to investigate NCV in the context of real-world, natural images, but also object-centric concept settings.
3.4 Verifiable Classification via the Merlin-Arthur Classifiers
@@
The second core component of NCV is a verifiable classifier of the Merlin-Arthur (MA) setup (Wäldchen et al.,, 2024) originally inspired by Interactive Proof Systems (Goldwasser et al.,, 1985). This setup generally formalizes the idea of proving that a classification decision is supported by a sparse and informative set of features.
@@
In NCV, specifically, two competing provers, Merlin and Morgana, select concept subsets either to support or mislead classification, respectively. The verifier, Arthur, then makes predictions based solely on these masked concepts (often denoted as “certificates” in PVG literature) without knowledge of the selector’s intent. Formally, each prover outputs a sparse binary mask with $m$ active entries, producing selected subsets:
$\mathcal{S}=M(\mathbf{c})\odot\mathbf{c},\quad\widehat{\mathcal{S}}=\widehat{M}(\mathbf{c})\odot\mathbf{c},$
where $\mathbf{c}$ is the concept encoding from extractor $g$, and $\odot$ denotes element-wise masking. Notably, all three agents represent differentiable neural models, with Arthur specifically representing a non-linear model.
This interactive setup enables two key metrics: (i) completeness, i.e., the probability that Arthur correctly classifies the input when using Merlin’s selected features and (ii) soundness, i.e., the probability that Arthur avoids incorrect classification when using Morgana’s features, either by giving the correct class or opting for a rejection. Overall, these components in NCV ensure that only sparse, faithful concept subsets are used and can be inspected for final predictions.

### v3

@@ Abstract
While Prover-Verifier Games (PVGs) offer a promising path toward verifiability in nonlinear classification models, they have not yet been applied to complex inputs such as high-dimensional images. Conversely, expressive concept encodings effectively allow to translate such data into interpretable concepts but are often utilised in the context of low-capacity linear predictors. In this work, we push towards real-world verifiability by combining the strengths of both approaches. We introduce Neural Concept Verifier (NCV), a unified framework combining PVGs for formal verifiability with concept encodings to handle complex, high-dimensional inputs in an interpretable way. NCV achieves this by utilizing recent minimally supervised concept discovery models to extract structured concept encodings from raw inputs. A prover then selects a subset of these encodings, which a verifier, implemented as a nonlinear predictor, uses exclusively for decision-making. Our evaluations show that NCV outperforms classic concept-based models and pixel-based PVG classifier baselines on high-dimensional, logically complex datasets and helps mitigate shortcut behavior. Overall, we demonstrate NCV as a promising step toward concept-level, verifiable AI.
@@ 1 Introduction
Deep learning has achieved remarkable predictive performances, but often at the expense of interpretability and trustworthiness (Rudin, 2019). However, particularly in high-stakes applications, it is critical that models provide verifiable justifications for their decisions (Irving et al., 2018; Fok and Weld, 2023). Prover-Verifier Games (PVGs), introduced by Anil et al. (2021), formalize such justifications via a game-theoretic approach: a prover provides evidence to convince a verifier, who accepts only verifiable proofs. A prominent instantiation of PVGs is the Merlin-Arthur Classifier (Wäldchen et al., 2024), a classifier guided by cooperative and adversarial provers, offering formal interpretability guarantees through information-theoretic bounds. Specifically, identified features are provably relevant for a given output, rather than mere correlations as in typical feature explanation methods (Adebayo et al., 2018; Zhou et al., 2022). However, the Merlin-Arthur framework faces significant scalability challenges when applied to high-dimensional real-world data, as explanations based on raw pixels are both computationally difficult to optimize and offer limited human understandability (cf. Fig.  1; Wäldchen et al. (2024)).
@@
Concurrently, Concept Bottleneck Models (CBMs) have emerged as a powerful framework for interpretable machine learning, structuring predictions through intermediate interpretable concept encodings (Koh et al., 2020; Stammer et al., 2021). Despite their advantages, CBMs typically employ linear classifiers on top of concept encoding layers, thereby potentially restricting expressivity and failing on tasks requiring nonlinear interactions among concepts (e.g., XOR problems, counting or permutation invariance, cf. Suppl.  B) (Mahinpei et al., 2021; Kimura et al., 2024; Lee et al., 2019).
@@
In this work, we combine the best of both worlds by introducing the Neural Concept Verifier (NCV), a novel framework integrating concept-based representations into PVGs in the form of Merlin-Arthur Classifiers. NCV shifts the prover–verifier interaction from the image level to a structured, symbolic concept level, overcoming both the scalability limitations encountered by Merlin-Arthur Classifiers in high-dimensional settings and the expressivity constraints inherent to linear CBMs. We demonstrate through extensive evaluations that NCV successfully scales PVGs to complex, high-dimensional classification tasks. At the same time, NCV enables verifiable, performant nonlinear classifiers on top of concept extractors, effectively narrowing the interpretability–accuracy gap in linear CBMs. Lastly, our framework improves robustness to shortcut learning, thereby enhancing the generalizability and trustworthiness of predictions, particularly in high-stakes applications.
@@
Our summarized contributions are: (i) We propose Neural Concept Verifier (NCV), a framework combining concept-based models with Prover–Verifier Games (PVGs). (ii) NCV scales PVGs to high-dimensional image data by operating on compact concept encodings. (iii) It enables expressive yet interpretable classification via sparse, nonlinear reasoning over concepts. (iv) We validate NCV on synthetic and real-world benchmarks, demonstrating strong accuracy and verifiability. (v) We highlight that NCV improves generalization under spurious correlations. The remainder of the paper is structured as follows. We begin with a review of related work, followed by an introduction and formal description of NCV. We then provide a comprehensive evaluation and, finally, discuss our findings and conclude the paper.
@@ 2 Background
Prover-Verifier Games. PVGs were introduced by Anil et al. (2021) as a game-theoretic framework to encourage learning agents to produce testable justifications through interactions between an untrusted prover and a trusted verifier. Their work showed that, under suitable conditions, the verifier can learn robust decision rules even when the prover actively attempts to persuade it of arbitrary outputs. Wäldchen et al. (2024) extended this idea to the Merlin-Arthur Classifier (MAC), which provides formal interpretability guarantees by bounding the mutual information between the selected features and the ground-truth label. Recently, Kirchner et al. (2024) applied a PVG-inspired approach to improve legibility of Large Language Model (LLM) outputs and Amit et al. (2024) introduced self-proving models that leverage interactive proofs to formally verify the correctness of model outputs. PVG-style setups have also been explored in safety-focused learning protocols (Irving et al., 2018; Brown-Cohen et al., 2024; Głuch et al., 2024). These developments reflect a broader trend of utilising multi-agent learning (Pruthi et al., 2022; Schneider and Vlachos, 2024; Du et al., 2024; Nair et al., 2023; Stammer et al., 2024a). Our work builds on this line of research by embedding PVGs into a concept-based classification framework to address the dimensionality bottleneck that limits their performance in high-dimensional settings.
@@
Concept Representations for Interpretability. The introduction of Concept Bottleneck Models (CBMs) (Koh et al., 2020; Espinosa Zarlenga et al., 2022) and concept-based explanations (Kim et al., 2018; Crabbé and van der Schaar, 2022; Poeta et al., 2023; Lee et al., 2025) was an important moment in the growing interest in AI interpretability research. The particular appeal of CBMs lies in the promise of interpretable predictions and a controllable, structured interface for human interactions (Stammer et al., 2021). While initial CBMs relied on fully supervised concept annotations, subsequent research has relaxed this requirement by leveraging pretrained vision-language models like CLIP for concept extraction (Bhalla et al., 2024; Yang et al., 2023; Oikarinen et al., 2023; Panousis et al., 2024; Steinmann et al., 2025), or employing fully unsupervised concept discovery methods (Ghorbani et al., 2019; Stammer et al., 2024b; Schut et al., 2025). While most works have focused on the bottleneck itself by reducing supervision requirements, mitigating concept leakage, or dynamically expanding the concept space, several recent approaches also enrich the classifier component using nonlinear or symbolic predictors, including concept-based memory reasoning (Debot et al., 2024), neural-symbolic reasoning (Barbiero et al., 2023), and causal concept models (Dominici et al., 2024; De Felice et al., 2025). NCV is complementary to these methods: rather than proposing a new classifier architecture, it wraps any concept-level predictor in a prover–verifier game that enforces sparse, per-sample concept selection and evaluates predictions under competing subsets.
Figure 2: Overview of the Neural Concept Verifier (NCV). The input image is first processed by a concept extractor to produce symbolic concept encodings. A prover–verifier game is then played over these encodings: a cooperative prover selects a sparse concept subset supporting the true class, while an adversarial prover selects misleading concepts. Finally, the nonlinear verifier makes a prediction based only on these selected concepts, ensuring verifiable and robust classification.
@@
Shortcut Learning. Independent of their interpretability, training deep models can often lead to unwanted artifacts and side effects. Shortcut learning describes the problem of models learning to rely on unwanted and unintended features to resolve a task (Geirhos et al., 2020). This is a common problem when training purely deep learning models (Lapuschkin et al., 2019; Schramowski et al., 2020) or neuro-symbolic models (Marconato et al., 2023; Bortolotti et al., 2025), and, if not taken care of, can lead to predictions being right for the wrong reasons (Ross et al., 2017). There have been various approaches to tackle this problem, from careful dataset curation (Ahmed et al., 2021) to modified model training (Friedrich et al., 2023), cf. Steinmann et al. (2024) for a comprehensive overview. While NCV is not specifically designed to mitigate shortcuts, we show that our setup can intrinsically mitigate their impact on the data.
@@ 3 Neural Concept Verifier (NCV)
In this section, we introduce the Neural Concept Verifier (NCV), a framework that combines concept-based representations with the Merlin-Arthur prover-verifier paradigm (Wäldchen et al., 2024). NCV trains nonlinear classifiers whose predictions provably rely on sparse subsets of high-level concepts, with information-theoretic guarantees formalized through completeness and soundness criteria (cf. Suppl.  A). In contrast to the original Merlin-Arthur classifier (MAC), which operates on raw pixel features and struggles to scale to high-dimensional inputs, NCV performs the prover–verifier interaction directly in concept space. This shift enables optimization on complex datasets and grounds PVG explanations in human-interpretable concepts.
NCV consists of two main components (cf. Fig.  2): (i) a weakly supervised concept extractor that transforms input data into interpretable concept encodings, and (ii) a nonlinear MAC that selects sparse concepts for the final predictions. After providing background notations and a high-level overview of NCV, we present detailed descriptions of each main component as well as training and inference details.
@@ 3.1 Problem Setup and Notation
[…]
@@ 3.2 The NCV Framework
NCV decomposes $f$ into the following components (cf. Fig.  2 for an illustrative overview):
@@
    A pair of prover agents $M,\widehat{M}:\mathcal{C}\rightarrow\{0,1\}^{C}$ that produce sparse binary masks selecting $m$ concepts each. $M$ (Merlin, cooperative prover) aims to help classification; $\widehat{M}$ (Morgana, adversarial prover) aims to mislead, which is crucial for overall robustness.
@@
    A nonlinear verifier (Arthur) $A:\mathbb{R}^{C}\rightarrow\mathcal{Y}$, which predicts a label based only on the masked concepts.
@@ 3.3 Concept Extraction
The concept extractor $g:\mathcal{X}\to\mathbb{R}^{C}$ transforms raw input data into interpretable, high-level concept representations, where each dimension corresponds to a semantically meaningful concept. The resulting concept encoding $\mathbf{c}\in\mathbb{R}^{C}$ serves as the input to the PVG.
While conceptually simple, a careful combination of PVGs and concept-based models is necessary. The concept extractor must satisfy three key requirements: interpretability, expressiveness, and modularity. Concepts should correspond to human-understandable features that can serve as meaningful explanations, the concept space should capture sufficient information for the target task without creating information bottlenecks, and the extractor should operate independently of the prover-verifier components. Unlike traditional concept extractor approaches that enforce sparsity constraints directly on $\mathbf{c}$, our framework delegates sparsity to the prover-verifier interaction. This allows the concept space to remain dense and expressive while achieving interpretable sparsity through downstream concept selection. Further, NCV can accommodate different supervision paradigms for concept extraction: supervised methods that leverage predefined concept vocabularies, self-supervised methods that exploit multi-modal correspondences (e.g., vision-language alignment), or unsupervised methods that discover latent conceptual structures can all be used as concept extractors.
Overall, NCV requires that $g$ produces consistent, interpretable encodings while maintaining sufficient information for accurate classification. In our evaluations, we instantiate NCV’s concept extractor via the recent unsupervised, object-centric NCB framework (Stammer et al., 2024b) and the multi-modal, CLIP-based SpLiCE (Bhalla et al., 2024). Operating in concept space rather than raw input space provides: (i) scalability through dimensionality reduction and (ii) explanations based on human-interpretable concepts.
3.4 Verifiable Classification via Merlin-Arthur Classifiers
@@
The second core component of NCV is a interpretable classifier of the Merlin-Arthur setup (Wäldchen et al., 2024) originally inspired by Interactive Proof Systems (Goldwasser et al., 1985). This setup generally formalizes the idea of proving that a classification decision is supported by a sparse and informative set of features.
@@
In NCV, specifically, two competing provers, Merlin and Morgana, select concept subsets either to support or mislead classification, respectively. The verifier, Arthur, then makes predictions based solely on these masked concepts without knowledge of the prover’s intent. Formally, each prover outputs a sparse binary mask with $m$ active entries, producing Merlin’s cooperative subset $S=M(\mathbf{c})\odot c$ and Morgana’s adversarial subset $\widehat{S}=\widehat{M}(\mathbf{c})\odot\mathbf{c}$, where $\mathbf{c}$ is the concept encoding from extractor $g$, and $\odot$ denotes element-wise masking. Notably, all three agents represent differentiable nonlinear models.
This interactive setup enables two key metrics: completeness and soundness. With Merlin’s subset $S$ and Morgana’s subset $\widehat{S}$, we define
Completeness $=\;\mathbb{P}_{(x,y)\sim\mathcal{D}}\!\big[A(S)=y\big],$ (1)
Soundness $=\;\mathbb{P}_{(x,y)\sim\mathcal{D}}\!\big[A(\widehat{S})\in\{y,\bot\}\big],$ (2)
where $\bot$ denotes the rejection class. Intuitively, completeness measures how often Arthur can recover the true label from Merlin’s sparse, helpful concepts (similar to standard accuracy), while soundness measures how often Arthur can avoid committing to a wrong label under Morgana’s misleading subset, either by staying correct or abstaining (cf. Suppl.  A for the theoretical interpretation).
Furthermore, sparsity has recently become central in concept-based models, as large concept spaces require sparse predictions for interpretability. Prior work (Bhalla et al., 2024; De Santis et al., 2025) achieves this by regularizing the concept space itself, restricting the number of active concepts before training the classifier. In contrast, NCV keeps the concept space fully expressive and enforces sparsity only in the concepts passed to the classifier.